a 40kg girl and an 8.4 kg sled are connected and are on frictionless ice of a frozen lake,15 m apart and the girl exerts 5.2N on the rope...how far from the girl's position do they meet?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

 

girl exerts 5.2N on the rope

so rope also exerts 5.2 N force on girl .and same force also exert ny rope at the other end to sled

acceleration of girl = 5.2/40  =.13 m/sec2

acceleration of sled = 5.2 /8.4 =.62 m/sec2

relative acceleration of girl w.r.t sld  =a = .62 + .13 =.75 m/sec2

let time taken to meet  is t

S=1/2 a t2

15 =1/2 * .75 * t2

 t2 =40

distance travel by girl in this time

S=1/2* 5.2/40 * 40

   =5.2/2 =2.6 m answer

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin